Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Weekly Remedial Activities Update
Report # 11: September 18 - September 22, 2006

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEK OF SEPTEMBER 18 - SEPTEMBER 22

  • Transportation and disposal of soils from stockpiles 001 – 003; approximately 200-cy remaining
  • Limited mechanical screening activities, screen off-line on 9/19/06
  • Perform re-excavation at the C.I.D.
  • Completed pre-characterization soil sampling in the Lagoon Area
  • Placed screened soils in the stockpile staging area for off-site transport
  • Completed pre-excavation, topographic survey of the East Fill Area
  • Continued to excavate the East Fill Area
  • Continued stockpile management activities, covering and securing tarpaulins
  • Consolidated oversize screened material
  • Collected confirmatory soil samples from excavations in the East Fill Area
  • Collected waste characterization soil samples from soils originating from East Fill excavation for disposal profiling for off-site landfill
  • Performed clearing of vegetation in Lagoon Area
  • Provided survey of post excavation limits in the C.I.D. and the East Fill Area
  • Continued to water the roads for dust control
  • Continue with air monitoring program

2. PROJECTED ACTIVITIES FOR THE WEEK OF SEPTEMBER 25 - SEPTEMBER 29

  • Stabilize failed TCLP Lead soils (pending TSP shipment)
  • Continue with East Fill excavation activities
  • Collect confirmatory soil samples from excavations in the East Fill Area
  • Collect waste characterization soil samples from soils originating from East Fill excavation for disposal profiling for off-site landfill
  • Re-place screen and continue to screen soils originating from East Fill Area
  • Place soils in the stockpile staging area
  • Collect additional clean backfill soil samples from the Lopes Construction fill source
  • Continue to water the haul roads for dust control
  • Continue to collect pre and post excavation survey data of the excavations
  • Continue water handling activities
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 14 September – 20 September 2006 can be reviewed in the table below.

Weekly Dust Monitoring Summary
Week #11: Results for Thursday through Wednesday**

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Hurricane Barrier (µg/m3) Church Street 2 (µg/m3)
9/14/2006 Precip Precip Precip Precip Precip
9/15/2006 Precip Precip Precip Precip Precip
9/18/2006 Precip Precip Precip Precip Precip
9/19/2006 45 20 0 NSC NSC
9/20/2006 10 40 0 1 0

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*: Monitoring conducted over 24-hour period.
**: No work and air monitoring occurs on weekend days

4. QUANTITY SUMMARY

Report # 11 - SUMMARY
TO DATE
ACTIVITIES TOTAL QUANTITY TOTAL QUANTITY
C.I.D. Excavation (Re-Digs) 100 CY 6,274 CY
East Fill Excavation 2,072 CY 3,160 CY
Soil Screening (CID) 0 CY 6,371 CY
Soil Screening (East Fill) 250 CY 750 CY
Soil placed in Stockpile Staging Area 300 CY *9,729 CY
Soils T&D to Crapo Landfill 437 Tons 1,238 Tons

Notes:
CY = volume of soil/debris in cubic yards
GAL = gallons of water

* Reported volume includes 2,655 CY of soils excavated from Commercial Area Excavation

Jump to main content.